OpenStudy (anonymous):

One number exceeds another by 16. The sum of the number is 58. What are the numbers?

OpenStudy (anonymous):

I set it up like this but I think its wrong x+16=58

OpenStudy (anonymous):

There are two unknown numbers x and y. Since the sum is 58 then x+y=58 one number is greater by 16 so that means y = x+16 substitute y into the first equation x+x+16=58 2x=42 x=21 y = 21+16 y=37 so your answer is 21 and 37

hero (hero):

x + y = 58 x = y + 16 y = 58 - x y = x - 16 y = y 58 - x = x - 16 x + x = 58 + 16 2x = 74 x = 74/2 x = 37 y = 21